# Improving GFlowNets for Text-to-Image Diffusion Alignment

### v1

@@ 2.2 GFlowNets
Furthermore, for any terminating state $\mathbf{x}$, we require $F(\mathbf{x})=R(\mathbf{x})$. In practice, these constraints can be transformed into tractable training objectives, as will be shown in Section 3. Based on GFlowNet theories in Bengio et al. (2023), if the DB criterion is satisfied for any transition, then the terminating distribution $P_{T}(\cdot)$ will be the same desired target distribution whose density is proportional to $R(\cdot)$.
3 Methodology
3.1 Denoising Markov decision process
The denoising process for text-to-image diffusion models can be easily reformulated as a multi-step Markov decision process (MDP) with finite horizon (Fan et al., 2023; Black et al., 2023) as follows:
@@ 3.1 Denoising Markov decision process
Training a standard RL algorithm within this diffusion MDP to perfection means the model would only generate a single trajectory with the largest reward value. This usually comes with the disadvantage of mode collapse in generated samples in practice. One direct solution is soft / maximum entropy RL (Ziebart et al., 2008; Fox et al., 2017; Haarnoja et al., 2017; Zhang et al., 2023c), whose optimal solution is a trajectory-level distribution and the probability of generating each trajectory is proportional to its trajectory cumulative reward, i.e., $p_{\boldsymbol{\theta}}(\mathbf{x}_{0:T})\propto\sum_{t}R_{t}(\mathbf{x}_{t})=R(\mathbf{x}_{0})$. However, in theory this means $p_{\boldsymbol{\theta}}(\mathbf{x}_{0})=\int p_{\boldsymbol{\theta}}(\mathbf{x}_{0:T})\mathop{}\!\mathrm{d}\mathbf{x}_{1:T}\propto\int R(\mathbf{x}_{0})\mathop{}\!\mathrm{d}\mathbf{x}_{1:T}=R(\mathbf{x}_{0})\cdot\int\mathop{}\!\mathrm{d}1\mathbf{x}_{1:T}$, which is not a well-defined finite term for unbounded continuous spaces. In contrast, the optimal solution of GFlowNet is $p_{\boldsymbol{\theta}}(\mathbf{x}_{0})\propto R(\mathbf{x}_{0})$.
@@
This formulation has a direct connection to the GFlowNet MDP definition in Section 2.2, which has been pointed out by Zhang et al. (2022a) and developed in Lahlou et al. (2023); Zhang et al. (2023b); Venkatraman* et al. (2024). To be specific, the action transition $(\mathbf{s}_{t},\mathbf{a}_{t})\to\mathbf{s}_{t+1}$ is a Dirac distribution and can be directly linked with the $(\mathbf{s}_{t}\to\mathbf{s}_{t+1})$ edge transition in the GFlowNet language. More importantly, the conditional distribution of the denoising process $p_{\boldsymbol{\theta}}(\mathbf{x}_{T-t-1}|\mathbf{x}_{T-t})$ corresponds to the GFlowNet forward policy $P_{F}(\mathbf{s}_{t+1}|\mathbf{s}_{t})$, while the conditional distribution of the diffusion process $q(\mathbf{x}_{T-t}|\mathbf{x}_{T-t-1})$ corresponds to the GFlowNet backward policy $P_{B}(\mathbf{s}_{t}|\mathbf{s}_{t+1})$. Besides, $\mathbf{x}_{t}$ is a GFlowNet terminal state if and only if $t=0$.
@@
3.2 Diffusion alignment with GFlowNets
In this section, we describe our proposed algorithm, diffusion alignment with GFlowNets (DAG). Rather than directly optimizing the reward targets as in RL, we aim to train the generative models so that in the end they could generate objects with a probability proportional to the reward function: $p_{\boldsymbol{\theta}}(\mathbf{x}_{0})\propto R(\mathbf{x}_{0})$. To achieve this, we construct the following DB-based training objective based on Equation 3, by regressing its one side to another in the logarithm scale for any diffusion step transition $(\mathbf{x}_{t},\mathbf{x}_{t-1})$.
$\ell_{\text{DB}}(\mathbf{x}_{t},\mathbf{x}_{t-1})=\left(\log F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)+\log p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},t)-\log F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},t-1)-\log q(\mathbf{x}_{t}|\mathbf{x}_{t-1})\right)^{2}$ (6)
We additionally force $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t=0)=R(\mathbf{x}_{0})$ to introduce the reward signal. Here ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ are the parameters of the diffusion U-Net model and the GFlowNet state flow function (which is another neural network), respectively. One can prove that if the optimization is perfect, the resulting model will generate a distribution whose density value is proportional to the reward function $R(\cdot)$ (Bengio et al., 2023; Zhang et al., 2023b).
Algorithm 1 Diffusion alignment with GFlowNets (DAG-DB & DAG-KL)
1:Denoising policy $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},t)$, noising policy $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$, flow function $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)$, black-box reward function $R(\cdot)$
2:repeat
3:     Rollout $\tau=\{\mathbf{x}_{t}\}_{t}$ with $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},t)$
4:     For each transition $(\mathbf{x}_{t},\mathbf{x}_{t-1})\in\tau$:
5:     if algorithm is DAG-DB then
6:          # normal DB-based update
7:         Update ${\boldsymbol{\theta}}$ and ${\boldsymbol{\phi}}$ with Equation 8
8:     else if algorithm is DAG-KL then
9:          # KL-based update
10:         Update ${\boldsymbol{\phi}}$ with Equation 8
11:         Update ${\boldsymbol{\theta}}$ with Equation 14
12:     end if
13:until some convergence condition
One way to parameterize the state flow function $F$ is through the so-called forward-looking (Pan et al., 2023b, FL) technique in the way of $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)=\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\mathbf{x}_{t})$, where $\tilde{F}_{\boldsymbol{\phi}}$ is the actual neural network to be learned. Intuitively, this is equivalent to initializing the state flow function to be the reward function in a functional way; therefore, learning of the state flow would become an easier task. Note that to ensure $F_{\boldsymbol{\phi}}(\mathbf{x}_{0},0)=R(\mathbf{x}_{0})$, we need to force $\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{0},0)=1$ for all $\mathbf{x}_{0}$ at the terminal step.
Incorporating denoising diffusion-specific structure
However, the intermediate state $\mathbf{x}_{t}$ is noisy under our context, and thus not appropriate for being evaluated by the given reward function, which would give noisy result. What’s more, what we are interested here is to “foresee” the reward of the terminal state $\mathbf{x}_{0}$ taken from the (partial) trajectory $\mathbf{x}_{t:0}$ starting from given $\mathbf{x}_{t}$. As a result, we can do the FL technique utilizing the particular structure of diffusion model as in $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)=\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t))$, where $\hat{\mathbf{x}}_{\boldsymbol{\theta}}$ is the data prediction network. We notice that a similar technique has been used to improve classifier guidance (Bansal et al., 2023). In short, our innovation in FL technique is
$F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)=\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\mathbf{x}_{t})\ \Longrightarrow\ F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)=\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)).$ (7)
Then the FL-DB training objective becomes
$\ell_{\text{FL}}(\mathbf{x}_{t},\mathbf{x}_{t-1})=\left(\log\frac{\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)R(\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t))p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{\tilde{F}_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})R(\hat{\mathbf{x}}_{\boldsymbol{\theta}}(\mathbf{x}_{t-1},{t-1}))q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}\right)^{2}.$ (8)
Remark 3 (GPU memory and the choice of GFlowNet objectives).
Similar to the temporal difference-$\lambda$ in RL (Sutton, 1988), it is possible to use multiple connected transition steps rather than a single transition step to construct the learning objective. Other GFlowNet objectives such as Malkin et al. (2022); Madan et al. (2022) use partial trajectories with a series of transition steps to construct the training loss and provide a different trade-off between variance and bias in credit assignment. However, for large-scale setups, this is not easy to implement, as computing policy probabilities for multiple transitions would correspondingly increase the GPU memory and computation multiple times. For example, in the Stable Diffusion setting, we could only use a batch size of $8$ on each GPU for single transition computation. If we want to use a two transition based training loss, we would need to decrease the batch size by half to $4$. Similarly, we will have to shorten the trajectory length by a large margin if we want to use trajectory balance. This may influence the image generation quality and also make it tricky to compare with the RL baseline, which can be implemented with single transitions and does not need to decrease batch size or increase gradient accumulation. In practice, we find that single transition algorithms (such as our RL baseline) perform reasonably well.
3.3 A KL-based GFlowNet algorithm with REINFORCE gradient
$\min_{\boldsymbol{\theta}}\mathcal{D}_{\mathrm{KL}}\left(p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\|\frac{F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}{F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)}\right).$ (9)
In theory, when DB is perfectly satisfied, the right term $F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})q(\mathbf{x}_{t}|\mathbf{x}_{t-1})/{F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)}$ is a normalized density; in practice, it could be an unnormalized one but does not affect the optimization. Next, define
$b(\mathbf{x}_{t},\mathbf{x}_{t-1})=\text{stop-gradient}\left(\log\frac{F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}\right),$ (10)
Proposition 4.
$\nabla_{\boldsymbol{\theta}}\mathcal{D}_{\mathrm{KL}}\left(p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\|\frac{F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},{t-1})q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}{F_{\boldsymbol{\phi}}(\mathbf{x}_{t},t)}\right)=\mathbb{E}_{\mathbf{x}_{t-1}\sim p_{\boldsymbol{\theta}}(\cdot|\mathbf{x}_{t})}\left[b(\mathbf{x}_{t},\mathbf{x}_{t-1})\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\right].$ (11)
We defer its proof to Section A.1 and make the following remarks:.
Remark 5 (gradient equivalence to detailed balance).
Remark 6 (analysis of $b(\mathbf{x}_{t},\mathbf{x}_{t-1})$).
The term $b(\mathbf{x}_{t},\mathbf{x}_{t-1})$ seems to serve as the traditional “reward” in the RL framework. Previous works (Tiapkin et al., 2024; Mohammadpour et al., 2024; Deleu et al., 2024) have shown that GFlowNet could be interpreted as solving a maximum entropy RL problem in a modified MDP, where any intermediate transition is modified to have an extra non-zero reward that equals the logarithm of GFlowNet backward policy: $r_{\text{mod}}(\mathbf{x}_{t},\mathbf{x}_{t-1})=\log q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),t>0$. What’s more, the logarithm of the transition flow function (i.e., either side of the detailed balance constraint of Equation 3) could be interpreted as the $Q$-function of this maximum entropy RL problem on the modified MDP. Therefore, we could take a more in-depth analysis of $-b(\mathbf{x}_{t},\mathbf{x}_{t-1})$:
$\underbrace{\log{F_{\boldsymbol{\phi}}(\mathbf{x}_{t-1},t-1)q(\mathbf{x}_{t}|\mathbf{x}_{t-1})}}_{\text{modified}\ Q-\text{function}}+\underbrace{\left(-\log p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\right)}_{\text{entropy}}-\underbrace{\log F_{\boldsymbol{\phi}}(\mathbf{x}_{t},{t})}_{\text{constant}}.$ (12)
The last term $F_{\boldsymbol{\phi}}(\mathbf{x}_{t},{t})$ is a constant for the $\mathbf{x}_{t-1}\sim p_{\boldsymbol{\theta}}(\cdot|\mathbf{x}_{t})$ process. Consequently, we can see that minimizing this KL is effectively equivalent to having a REINFORCE gradient with the modified $Q$-function plus an entropy regularization (Haarnoja et al., 2018). What’s more, the entropy of forward policy $\mathbb{E}_{\mathbf{x}_{t-1}\sim p_{\boldsymbol{\theta}}(\cdot|\mathbf{x}_{t})}\left[-\log p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\right]$ is actually also a constant (the diffusion model’s noise schedule is fixed), which does not affect the learning.
Therefore, we can define a new objective to be
$\ell_{\text{KL}}(\mathbf{x}_{t},\mathbf{x}_{t-1})=b(\mathbf{x}_{t},\mathbf{x}_{t-1})\ \text{clip}\left(\frac{p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{p_{{\boldsymbol{\theta}}_{\text{old}}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})},1-\epsilon,1+\epsilon\right),$ (14)
where $\mathbf{x}_{t-1}\sim p_{{\boldsymbol{\theta}}_{\text{old}}}(\cdot|\mathbf{x}_{t})$. Here we also introduce a clip operation to remove too drastic update, following PPO (Schulman et al., 2017). In this way, the overall gradient along a trajectory becomes
$\mathbb{E}_{p_{{\boldsymbol{\theta}}_{\text{old}}}(\mathbf{x}_{0:T})}\left[\sum_{t=1}^{T}b(\mathbf{x}_{t},\mathbf{x}_{t-1})\nabla_{\boldsymbol{\theta}}\ \text{clip}\left(\frac{p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})}{p_{{\boldsymbol{\theta}}_{\text{old}}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})},1-\epsilon,1+\epsilon\right)\right].$ (15)
We use this to update the policy parameter ${\boldsymbol{\theta}}$ and use FL-DB to only update ${\boldsymbol{\phi}}$. We call this “diffusion alignment with GFlowNet and REINFORCE gradient” method to be DAG-KL. Note that when calculating $b(\mathbf{x}_{t},\mathbf{x}_{t-1})$, we also adopt the diffusion-specific FL technique developed in Section 3.2. We also put the algorithmic pipeline of DAG-KL in Algorithm 1.
@@ 3.3 A KL-based GFlowNet algorithm with REINFORCE gradient
4 Related Works
Diffusion alignment
People have been modeling human values to a reward function in areas such as game (Ibarz et al., 2018) and language modeling (Bai et al., 2022) to make the model more aligned. In diffusion models, early researchers used various kinds of guidance (Dhariwal & Nichol, 2021; Ho & Salimans, 2022) to achieve the goal of steerable generation under the reward. This approach is as simple as plug-and-play but requires querying the reward function during inference time. Another way is to post-train the model to incorporate the information from the reward function, which has a different setup from guidance methods; there is also work showing that this outperforms guidance methods (Uehara et al., 2024). Lee et al. (2023); Dong et al. (2023) achieve this through maximum likelihood estimation on model-generated samples, which are reweighted by the reward function. These works could be thought of as doing RL in one-step MDPs. Black et al. (2023); Fan et al. (2023) design RL algorithm by taking the diffusion generation process as a MDP (Section 3.1). In this work, we focus on black-box rewards where it is appropriate to use RL or GFlowNet methods. Furthermore, there are methods developed specifically for differentiable rewards setting (Clark et al., 2023; Wallace et al., 2023b; Prabhudesai et al., 2023; Wu et al., 2024; Xu et al., 2023; Uehara et al., 2024; Marion et al., 2024). Besides, Chen et al. (2023) study the effect of finetuning text encoder rather than diffusion U-Net. There is also work that relies on preference data rather than an explicit reward function (Wallace et al., 2023a; Yang et al., 2024). Kim et al. (2024a) investigate how to obtain a robust reward based on multiple different reward functions.
GFlowNets
GFlowNet is a family of generalized variational inference algorithms that treats the data sampling process as a sequential decision-making one. It is useful for generating diverse and high-quality samples in structured scientific domains (Jain et al., 2022; 2023b; Liu et al., 2022; Jain et al., 2023a; Shen et al., 2023; Zhang et al., 2023e; Pan et al., 2023a; Kim et al., 2023; 2024b). A series of works have studied the connection between GFlowNets and probabilistic modeling methods (Zhang et al., 2022b; Zimmermann et al., 2022; Malkin et al., 2023; Zhang et al., 2022a; Ma et al., ), and between GFlowNets and control methods (Pan et al., 2023c; d; b; Tiapkin et al., 2024). GFlowNets also have wide application in causal discovery Deleu et al. (2022), phylogenetic inference (Zhou et al., 2024), and combinatorial optimization (Zhang et al., 2023a; d). A concurrent work (Venkatraman* et al., 2024) also studies GFlowNet on diffusion alignment which is similar to this work but has different scope and different developed algorithm. Specifically, this work is aiming for posterior approximate inference that the reward function is treated as likelihood information, and develops a trajectory balance (Malkin et al., 2022) based algorithm on length modified trajectories.
5 Experiments
Experimental setups
We choose Stable Diffusion v1.5 (Rombach et al., 2021) as our base generative model. For training, we use low-rank adaptation (Hu et al., 2021, LoRA) for parameter efficient computation. As for the reward functions, we do experiments with the LAION Aesthetics predictor, a neural aesthetic score trained from human feedback to give an input image an aesthetic rating. For text-image alignment rewards, we choose ImageReward (Xu et al., 2023) and human preference score (HPSv2) (Wu et al., 2023). They are both CLIP (Radford et al., 2021)-type models, taking a text-image pair as input and output a scalar score about to what extent the image follows the text description. We also test with the (in)compressibility reward, which computes the file size if the input image is stored in hardware storage. As for the prompt distribution, we use a set of $45$ simple animal prompts from Black et al. (2023) for the Aesthetics task; we use the whole imagenet classes for the (in)compressibility task; we use the DrawBench (Saharia et al., 2022) prompt set for the ImageReward task; we use the photo and painting prompts from the human preference dataset (HPDv2) (Wu et al., 2023) for the HPSv2 task. We notice that in our experiments, we use prompt set containing hundreds of prompts which is more than some previous work such as Black et al. (2023).
@@ 5 Experiments
Effectiveness of the proposed methods
We first demonstrate that our proposed methods could generated images that have meaningful improvements corresponding to the rewards being used. In Figure 1, we compare the images from the original Stable Diffusion pretrained model and our proposed method. After our post-training, the generated images become more vibrant and vivid; we also notice that these images have slightly higher saturation, which we believe is aligned with the human preference on good-looking pictures. We also visualize the experiment results on compressibility and incompressibility tasks in Figure 2. The second row shows the generated images from the model trained with the compressibility reward, which have low details and smooth textures, and also have very limited colors. On the other hand, the model trained with incompressibility reward would generate images with high frequency texture, as shown in the third row. These results indicate that our method could effectively incorporate the reward characteristics into the generative models. We defer more experimental details to Section B.1.
Algorithmic comparisons
The main baseline we compare with is denoising diffusion policy optimization (Black et al., 2023, DDPO), an RL algorithm that is specifically designed for denoising diffusion alignment and has been shown to outperform other align-from-black-box-reward methods including (Lee et al., 2023; Fan et al., 2023). We show the reward curves w.r.t. the training steps of the aesthetic, ImageReward, and HPSv2 rewards in Figure 3. Here, the number of training steps corresponds proportionally to the number of trajectories collected. Both our proposed methods, DAG-DB and DAG-KL, achieve faster credit assignment than the DDPO baseline by a large margin. We also put corresponding curve plots for compressibility and incompressibility rewards in Figure 4, which also demonstrates the advantage of our methods. We defer related training details to Section B.1.
@@
Appendix A Proof
A.1 Proof of Proposition 4
Proof.
Appendix B More about experiments
B.1 Experimental details

### v2

@@ A.2 GFlowNets
Furthermore, for any terminating state $\mathbf{x}$, we require $F(\mathbf{x})=R(\mathbf{x})$. In practice, these constraints can be transformed into tractable training objectives, as will be shown in Section 2. Based on GFlowNet theories in bengio2021foundations, if the DB criterion is satisfied for any transition, then the terminating distribution $P_{T}(\cdot)$ will be the same desired target distribution whose density is proportional to $R(\cdot)$.
Appendix B Related Works
Diffusion alignment
People have been modeling human values to a reward function in areas such as game (ibarz2018reward) and language modeling (bai2022training) to make the model more aligned. In diffusion models, early researchers used various kinds of guidance (dhariwal2021diffusion; ho2022classifierfree) to achieve the goal of steerable generation under the reward. This approach is as simple as plug-and-play but requires querying the reward function during inference time. Another way is to post-train the model to incorporate the information from the reward function, which has a different setup from guidance methods; there is also work showing that this outperforms guidance methods (uehara2024finetuning). Lee2023AligningTM; dong2023raft achieve this through maximum likelihood estimation on model-generated samples, which are reweighted by the reward function. These works could be thought of as doing RL in one-step MDPs. Black2023TrainingDM; Fan2023DPOKRL design RL algorithm by taking the diffusion generation process as a MDP (Section D.1). In this work, we focus on black-box rewards where it is appropriate to use RL or GFlowNet methods. Furthermore, there are methods developed specifically for differentiable rewards setting (clark2023directly; Wallace2023EndtoEndDL; prabhudesai2023aligning; wu2024deep; Xu2023ImageRewardLA; uehara2024finetuning; marion2024implicit). Besides, Chen2023EnhancingDM study the effect of finetuning text encoder rather than diffusion U-Net. There is also work that relies on preference data rather than an explicit reward function (wallace2023diffusion; yang2024using). kim2024confidenceaware investigate how to obtain a robust reward based on multiple different reward functions.
GFlowNets
GFlowNet is a family of generalized variational inference algorithms that treats the data sampling process as a sequential decision-making one. It is useful for generating diverse and high-quality samples in structured scientific domains (jain2022biological; jain2022multiobjective; Liu2022GFlowOutDW; Jain2023GFlowNetsFA; Shen2023TowardsUA; zhang2023distributional; pan2023pretraining; Kim2023LocalSG; kim2024learning). A series of works have studied the connection between GFlowNets and probabilistic modeling methods (zhang2022generative; Zimmermann2022AVP; malkin2022gfnhvi; zhang2022unifying; MaBakingSI), and between GFlowNets and control methods (pan2022gafn; pan2023stochastic; pan2023better; tiapkin2024generative). GFlowNets also have wide application in causal discovery deleu2022bayesian, phylogenetic inference (zhou2024phylogfn), and combinatorial optimization (Zhang2023RobustSW; Zhang2023LetTF). A concurrent work (gfninverse2024) also studies GFlowNet on diffusion alignment which is similar to this work but has different scope and different developed algorithm. Specifically, this work is aiming for posterior approximate inference that the reward function is treated as likelihood information, and develops a trajectory balance (malkin2022trajectory) based algorithm on length modified trajectories.
Appendix C Proof
C.1 Proof of Proposition 2
Proof.
Appendix D More about methodology
D.1 Denoising Markov decision process
The denoising process for text-to-image diffusion models can be easily reformulated as a multi-step Markov decision process (MDP) with finite horizon (Fan2023DPOKRL; Black2023TrainingDM) as follows:
@@ D.1 Denoising Markov decision process
Training a standard RL algorithm within this diffusion MDP to perfection means the model would only generate a single trajectory with the largest reward value. This usually comes with the disadvantage of mode collapse in generated samples in practice. One direct solution is soft / maximum entropy RL (ziebart2008maximum; fox2017taming; haarnoja2017reinforcement; zhang2023latent), whose optimal solution is a trajectory-level distribution and the probability of generating each trajectory is proportional to its trajectory cumulative reward, i.e., $p_{\boldsymbol{\theta}}(\mathbf{x}_{0:T})\propto\sum_{t}R_{t}(\mathbf{x}_{t})=R(\mathbf{x}_{0})$. However, in theory this means $p_{\boldsymbol{\theta}}(\mathbf{x}_{0})=\int p_{\boldsymbol{\theta}}(\mathbf{x}_{0:T})\mathop{}\!\mathrm{d}\mathbf{x}_{1:T}\propto\int R(\mathbf{x}_{0})\mathop{}\!\mathrm{d}\mathbf{x}_{1:T}=R(\mathbf{x}_{0})\cdot\int\mathop{}\!\mathrm{d}1\mathbf{x}_{1:T}$, which is not a well-defined finite term for unbounded continuous spaces. In contrast, the optimal solution of GFlowNet is $p_{\boldsymbol{\theta}}(\mathbf{x}_{0})\propto R(\mathbf{x}_{0})$.
@@
This formulation has a direct connection to the GFlowNet MDP definition in Section A.2, which has been pointed out by zhang2022unifying and developed in lahlou2023cgfn; Zhang2023DiffusionGF; gfninverse2024. To be specific, the action transition $(\mathbf{s}_{t},\mathbf{a}_{t})\to\mathbf{s}_{t+1}$ is a Dirac distribution and can be directly linked with the $(\mathbf{s}_{t}\to\mathbf{s}_{t+1})$ edge transition in the GFlowNet language. More importantly, the conditional distribution of the denoising process $p_{\boldsymbol{\theta}}(\mathbf{x}_{T-t-1}|\mathbf{x}_{T-t})$ corresponds to the GFlowNet forward policy $P_{F}(\mathbf{s}_{t+1}|\mathbf{s}_{t})$, while the conditional distribution of the diffusion process $q(\mathbf{x}_{T-t}|\mathbf{x}_{T-t-1})$ corresponds to the GFlowNet backward policy $P_{B}(\mathbf{s}_{t}|\mathbf{s}_{t+1})$. Besides, $\mathbf{x}_{t}$ is a GFlowNet terminal state if and only if $t=0$.
@@
D.2 Experimental details